OpenStudy (anonymous):

The data represent the number of phone calls received over a 7-day period. Day 1 2 3 4 5 6 7 Calls Received 5 2 9 8 3 1 0 What is the mean number of phone calls received? i think 4

OpenStudy (anonymous):

@pinklion23

OpenStudy (pinklion23):

im hear

OpenStudy (anonymous):

u mean here

OpenStudy (pinklion23):

okay so your number set is 5 2 9 8 3 1 0

OpenStudy (pinklion23):

now add them all up

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i did, i just want to make sure i am correct

OpenStudy (anonymous):

in the question

OpenStudy (pinklion23):

okay well you are good job lol :)

OpenStudy (texaschic101):

28/7 = 4...correct

OpenStudy (firejay5):

@Sworewolf You were correct the answer is 4, because you would add 5 + 2 + 9 + 8 + 3 + 1 + 0 = 28, then 28/7 = 4